OpenStudy (brandon222):

Which of the following sequences of numbers are arithmetic sequences? Check all that apply. A. -4, -7, -10, -13, -16, ... B. 8, 10.5, 13, 15.5, 18, ... C. 3, -3, 3, -3, 3, ... D. 1, 5, 10, 15, 20, ... E. 1000, 1001, 1002, 1003, 1004, ...

OpenStudy (welshfella):

an arithmetic sequence increases by a constant amount like 2, 5 , 8 , 11..... -this one has a a common difference of 3

OpenStudy (brandon222):

So it would be a b d e

OpenStudy (welshfella):

- note the common difference may also be negative for example -9,-11,-12,-13 here common difference is -1

OpenStudy (welshfella):

No there are only 2 correct choices

OpenStudy (brandon222):

D E

OpenStudy (welshfella):

E is correct because c d = 1 re check d - Note that the difference must be CONSTANT

OpenStudy (welshfella):

5 - 1 = 4 10 - 5 = 5 - the difference is not constant.

OpenStudy (welshfella):

so nits not d